OpenStudy (anonymous):

MEDAL WILL BE AWARDED!! Solve the quadratic equation by completing the square. -3x^2+9x=1

OpenStudy (anonymous):

solve -3x2 + 9x = 1 ⇒ x=0.1156 , x=2.884

OpenStudy (anonymous):

oh -_-

OpenStudy (anonymous):

How did you do that? Even if it's not completing the square...

OpenStudy (anonymous):

Uh...maybe igreen can help.....

OpenStudy (anonymous):

Call his number

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Sorry..I hate 'Completing the Square' I never understood it.

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Any other idea how to get the answer?

OpenStudy (anonymous):

I can send you some links ^_^

OpenStudy (anonymous):

you im only in 6th grade and watching all these problems that look like scrambled numbers

OpenStudy (anonymous):

You are little too young for this site

OpenStudy (anonymous):

how

OpenStudy (anonymous):

You can't really help people, most kids are teens

OpenStudy (anonymous):

@!Call~Me~Addi! How did you get that? @PaulaLovesSchool13 Please do.

OpenStudy (anonymous):

hmm ok

OpenStudy (anonymous):

ok I am searching

OpenStudy (alekos):

-3x^2 + 9x=1 -3x^2 + 9x -1 =0 3x^2 -9x +1 =0 x^2 - 3x +1/3 =0

OpenStudy (anonymous):

I agree @alekos

OpenStudy (anonymous):

http://www.purplemath.com/modules/solvquad.htm

OpenStudy (alekos):

(x^2 - 3x + 9/4) -9/4 +1/3 =0

OpenStudy (alekos):

(x-3/2)^2 -9/4 +1/3 =0

OpenStudy (alekos):

(x-3/2)^2 - 27/12 + 4/12 = 0

OpenStudy (alekos):

(x-3/2)^2 - 23/12 = 0 (x-3/2)^2 = 23/12 x-3/2 = sqrt(23/12) x = sqrt(23/12) + 3/2

OpenStudy (alekos):

did you follow all of that?

OpenStudy (alekos):

x is also equal to -sqrt(23/12) + 3/2

OpenStudy (anonymous):

I think so...

OpenStudy (alekos):

just go through all my steps one by one and you'll see how I've come to the final two answers

OpenStudy (anonymous):

He is confused

OpenStudy (alekos):

so x = 2.8844 and 0.1156

OpenStudy (alekos):

derived by completing the square

OpenStudy (anonymous):

So this was complete the square?